OpenStudy (anonymous):

Find an equation of a line that has the following: Slope=2, passes through (-5,-10) Slope= -1/3, passes through (-9,8)

OpenStudy (wilder.monday):

1: y + 10 = 2(x + 5)

OpenStudy (wilder.monday):

then simplify

OpenStudy (wilder.monday):

2: y - 8 = -1/3(x + 9). then simplify

OpenStudy (wilder.monday):

@pug